OpenStudy (anonymous):

is 8cos(3x-pi/3) if x=pi/6 exact answer -4squareroot3? (sorry idk how to make the symbols

OpenStudy (mertsj):

Well if x = pi/6 then 3x is pi/2 pi/2-pi/3=pi/6 cos(pi/6)= sqrt3/2 8cos (pi/6)=4sqrt3